OpenStudy (anonymous):

Input the equation of the given line. The line including (4, 7) and (0, 7)?

OpenStudy (anonymous):

y-y1=m(x-x1) find the slope slope=(y2-y1)/(x2-1)

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

It would help if I didn't miss the basic steps, huh?

OpenStudy (anonymous):

well do you want me to show or you would want to try yourself

OpenStudy (anonymous):

It would definitely be more helpful if you showed me!

OpenStudy (anonymous):

slope=0,the points lie on the line y=7

OpenStudy (anonymous):

Ah. Thank you so much!